### v1

@@ 8 Concluding remarks
In this paper, we have presented sharp estimates for the Schur complements arising in saddle point problems, by utilizing the auxiliary space theory introduced in [33]. We also provided various examples, such as the augmented Lagrangian method, mixed finite element methods, and nonoverlapping domain decomposition methods, which highlight the usefulness of our results. In particular, the proposed framework simplifies the analysis compared to existing approaches by minimizing the algebraic effort required. We believe that the framework not only offers deeper insight into saddle point problems but also serves as a valuable tool for the design and analysis of new efficient iterative methods.
@@
The second remark concerns the extension of the proposed framework to a broader range of problems. As mentioned at several points in this paper (see, e.g., Remark 3.4), some parts of the contents of this paper can already be extended to more general classes of convex optimization problems, although a full generalization is highly nontrivial. We note that there has recently been significant research on extending the theory of iterative methods for linear systems to convex optimization problems; see [14] and the references therein. In this spirit, we view the extension of our results to convex optimization problems as an important direction for future work, which we expect to be highly impactful.

### v2

@@ 8 Concluding remarks
In this paper, we have presented sharp estimates for the Schur complements arising in saddle point problems, by utilizing the auxiliary space theory introduced in [35]. We have also provided various examples, such as the augmented Lagrangian method, mixed finite element methods, and nonoverlapping domain decomposition methods, which highlight the usefulness of our results. In particular, the proposed framework simplifies the analysis compared to existing approaches by minimizing the algebraic effort required. We believe that the framework not only offers deeper insight into saddle point problems but also serves as a valuable tool for the design and analysis of new efficient iterative methods.
@@
The second remark concerns the extension of the proposed framework to a broader range of problems. As mentioned at several points in this paper (see, e.g., Remark 3.4), some of the results in this paper can already be extended to more general classes of convex optimization problems, although a full generalization is highly nontrivial. We note that there has recently been significant research on extending the theory of iterative methods for linear systems to convex optimization problems; see [16] and the references therein. In this spirit, we view the extension of our results to convex optimization problems as an important direction for future work, which we expect to be highly impactful.